So I'm having to simplify the square root of this problem with "x" and "y" having exponents. Afterward I check my answer by raising it to the 2nd power. I have looked through my text and online at examples but I feel like I'm missing something. The problem is:
sqrt64x^14y^20
Simplified, it is 8x^14y^20. Once I raise it to the second power, it changes to 64x^28y^40
I feel like the power of "x" and "y" should have changed when simplified. Should it?

Answer by rothauserc(4718)

if you mean
:
square root(64 * x^14 * y^20) where * means multiply, then
:
square root(64 * x^14 * y^20) = + or - (8 * x^7 * y^10)
:
when like bases are multiplied, you add the exponents, i.e.
:
x^14 = x^7 * x^7